A drama class has a total of 44 students. The number of males is 8 more than the number of females. How many males and how many females are in the class?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52810)

.

Let M be the number of males 
and F be the number of females.


Write two equations as you read the problem

    M + F = 44    (1)    (a total of 44 students)

    M - F =  8    (2)    (The number of males is 8 more than the number of females)


Add equations (1) and (2)  (both sides). You will get

    2M = 44 + 8 = 52.


Hence,  M = 52/2 = 26.

Then from equation (1),  F = 44-26 = 18.


?ANSWER.  26 males, 18 females.

Other solution methods are applicable and possible, too,
but this one is the most educative (in my point of view).

----------------

If you want to see many other similar (and different) solved problems of this kind with explanations,
look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.

Learn the subject from there and become a best student in the class / (in the school) / (in the district).



Answer by greenestamps(13200)



Let the number of females be x.

Then, since the number of males is 8 more than the number of females, the number of males is x+8.,br>
The total number of students is 44:

x+(x+8) = 44
2x+8 = 44
2x = 36
x = 18

ANSWERS:
number of females: x = 18
number of males: x+8 = 26